OpenStudy (anonymous):

pleeeease help! there are 7 teams (A,B,C,D,E,F,G) participating in a swimming tournament; each team 4 members. Question: 10 members are to be randomly selected from the 28 members. In now many ways can the selection process be conducted if there is a ruling that at most 2 members from the same team may be selected?

OpenStudy (alexwee123):

actually i'm 100% i did something wrong :/

OpenStudy (anonymous):

is it 161

OpenStudy (paxpolaris):

If you were to pick 1 person from each team ... you would have to pick a 2nd person from 3 of the teams

OpenStudy (paxpolaris):

Now, you could choose 2 people each from five of the teams: (7 choose 5) ways to pick 5 teams in 1 team (4choose2) ways to pic 2 people (4 choose 2)^5 ways to pick 2 people each from the 5 chosen teams So, number of ways you could choose: 2 people each, from 5 of the 7 teams\[\Large={7\choose5}\times{4\choose2}^5\]\[=21\times6^5\]

OpenStudy (paxpolaris):

OR, you could choose: ..... 2 people each from 4 of the teams, and ..... 1 person each from 2 of the remaining teams. \[\Large = {7 \choose 4}{4\choose2}^4\times {3\choose 2}{4\choose1}^2\]\[=35\cdot6^4\times3\cdot4^2\]

OpenStudy (paxpolaris):

OR, you could choose: ..... 2 people each from 3 of the teams, and ..... 1 person each from 4(all) of the remaining teams. \[\Large ={7\choose3}{4\choose2}^3\times \cancel{4\choose4} {4\choose1}^4\]\[=35\cdot6^3\times 1\cdot 4^4\]

OpenStudy (barrycarter):

The answer involves the hypergeometric distribution, but let's try to find it using recursion. How can you construct a solution for 7 teams and 10 members?: - Any solution for 6 teams and 10 members still works. - If you have a solution for 6 teams and 9 members, you can choose 1 member from team 7 (there are four ways to do this). - If you have a solution for 6 teams and 8 members, you can choose 2 members from team 7 (there are C[4,2] or 6 ways to do this). In other words, f[7 teams, 10 members] = f[6 teams, 10 members] + 4*f[6 teams, 9 members] + 6*f[6 teams, 8 members] More generically: f[t,m] = f[t-1,m] + 4*f[t-1,m-1] + 6*f[t-1,m-2] Our base case is 1 team. We can choose 1 member 4 ways and 2 members 6 ways. We can't choose any teams with 3 or 4 members, since this would break the rules. And, of course, we can't choose more than 4 members from a 4 member team. f[1,1] = 4 f[1,2] = 6 f[1,3] = 0 f[1,4] = 0 f[1,x] = 0 for x>=5 The answer I got after slogging through all the recursion is 3,918,240. There is almost definitely a better way to do this.

OpenStudy (paxpolaris):

@barrycarter ... i think maybe you forgot to include f[1,0] = 1 f[2,0] = 1 I really think my answer of 4,275,936 is correct

OpenStudy (barrycarter):

You're right, I'm wrong. Rats! Adding those two base cases gives me your answer.